# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ 4.2.2 Derivation without heuristics
In reading this figure, it must be noted that “aging-/slip-like” (i.e. $|f-f_{t}|\lessgtr|f-f_{\delta}V|$) evolution is distinct from evolution “conforming the aging/slip canon” (i.e. $f\simeq f_{t}$, $f\simeq f_{\delta}V$). Rigorous confirmation of the VS-canon conformity of laboratory data is limited to VS tests with $V_{\rm after}/V_{\rm before}>10^{-2}$. VS tests with further lower $V_{\rm after}/V_{\rm before}$ (down to $10^{-3.5}$ so far) are slip-like but can significantly differ from the VS canon (Bhattacharya et al. 2015; 2022). Figure 9 leaves this range unpainted. On the other hand, as for SHS, rigorous studies do not exist, to our knowledge, on how closely the laboratory data obeys the aging canon. There is even a possibility that (some of) the results by Nakatani & Mochizuki (1996) are just aging-like, not canon-conforming. Under those circumstances, while we painted the whole $(V,\theta)$ space of Nakatani & Mochizuki (1996) down to $\Delta\tau/A=8$ as “canon-conforming,” the formal conclusion of the present paper remains $\beta\lesssim 0.01$, with the lower bound for $\beta$ unspecified (§4.1). Nevertheless, finite $\beta$ is instantly concluded mathematically from the SHS tests under $-\log_{10}\Omega\gg 1$ and the VS behavior under $|\log_{10}\Omega|=\mathcal{O}(1)$, not suffering from the current inability to determine the $\beta$ value. Figure 9, though drawn in terms of the concrete values of $\beta$ inferred from experiments probing the currently limited range of $(V,\theta)$ space, is presented only to help readers grasp the general picture of the presently proposed aging-slip switch (i.e., the existence of $\beta$ for $f(V,\theta)$), which has been mathematically proven in §4.2 by the deduction from the canons (plus the premise of the TDF).
@@ 4.3 Examples of evolution laws consistent with the three canonical behaviors
In §4.2, we have shown that a class of evolution laws satisfying eqs.(80) and (83) conform constraints from SS, VS, and SHS experiments. The idea is a simple reconciliation of the slip and aging laws (Kato & Tullis, 2001; Ampuero & Rubin, 2008) by switching the two laws according to the value of $\Omega$ ($:=V\theta/L$), the conventional state variable $\theta$ normalized by its steady state value for the current slip rate $\theta_{\rm SS}(V)$ [$\Omega=\theta/\theta_{\rm SS}(V)$]. The proposed class of laws approximate $\dot{\theta}\simeq 1$ (the aging law at the NSC) for $\Omega\ll\beta$, while they approximate the canonical evolution in VS tests, $\dot{\theta}\simeq-(L/D_{\rm c})\Omega\ln\Omega$ (the slip law). In this subsection §4.3, we propose concrete examples of this class of evolution laws. A stepwise function $\psi_{\rm c}=H(\beta-\Omega)$, where $H(\cdot)$ is the Heaviside function, satisfies eq. (83), of course, but it causes an artificial discontinuity in the evolution function. Instead, we seek a procedure to construct smooth functions satisfying eq. (83). From this subsection onward, we assume $C_{\rm SS}=1$ (eq. 76), whereas §4.2 was developed under a less restrictive assumption of $C_{\rm SS}\sim 1$. In addition, within this subsection §4.3, we impose $L\sim D_{\rm c}$ (eq. 70) to simplify the derivation. As discussed in §3.4, $L\sim D_{\rm c}$ is likely to be the case, if not well confirmed by a wide range of experiments.
Constructing $\psi_{\rm c}$ for $\Omega\gg\beta$ so that it satisfies the first line of eq. (83) is not as straightforward as it might sound. The difficulty comes from $|f_{t|{\rm VS}}|\ll|f_{\delta|{\rm VS}}|V$ (eq. 44) because $f_{\delta}V$ ($=-(L/D_{\rm c})\Omega\ln\Omega$) can become infinitesimally small in the vicinity of $\Omega=1$. To get around, we first make a $\psi_{\rm c}$ that conforms eq. (83) except in the vicinity of $\Omega=1$. Exception handling is next made so that eq. (83) is satisfied also in the vicinity of $\Omega=1$. For this purpose, we introduce here a positive constant $\epsilon_{\Omega}^{*}$ and divide the $\Omega$ range into $|\Omega-1|<\epsilon_{\Omega}^{*}$ and the rest (Fig. 10). To simplify inequality evaluations in the rest of the derivation, we assume $1-\epsilon_{\Omega}^{*}>1/e$, or equivalently $\epsilon_{\Omega}^{*}<1-1/e=0.6321...$, noting that $\Omega\ln\Omega$ takes the unique extremum value at $\Omega=1/e$ (Fig. 10). Also, we hereafter assume $\beta<1/e^{2}(=0.135...)$ and $\sqrt{\beta}\ll 1$, again for simplicity of derivation. As discussed in §4.2, observed symmetry in many existing VS tests has already rejected the possibility that $\beta\geq 1/e^{2}$. Moreover, recent VS tests employing very large steps by Bhattacharya et al. (2015) show that $\sqrt{\beta}\lesssim 0.1$ (eq. 85).
@@
which follows eq. (80) and eq. (76) ($\theta_{\rm SS}=L/V$). In the first place, eq. (34) derived from the VS requirement intends that the slip-driven evolution alone makes $\Omega$ approach exactly to its stationary point $\Omega=1$. Hence, involvement of any purely time-dependent healing within the neighborhood of $\Omega=1$, even minuscule, would shift the steady state from $\Omega=1$, meaning that $C_{\rm SS}$ would deviate from unity. Therefore, $\psi_{\rm c}$ needs to be truncated for $\Omega$ almost at the steady state. A straightforward way to realize eq. (91) is to impose $\psi_{\rm c}=0$ in $\Omega>1-\epsilon_{\Omega}^{*}$:
$\psi_{\rm c}(\Omega):=\psi_{\rm cBase}(\Omega)H(1-\epsilon^{*}_{\Omega}-\Omega).$ (93)
@@
Hereafter, we refer to eq. (96) as the ‘modified composite law’ because its form under $L=D_{\rm c}$ is exactly the composite law (eq. 74) by Kato & Tullis (2001), with the constant $V_{\rm c}$ in eq. (74) replaced by $\beta L/\theta$. To sum up, by turning on/off the purely time-dependent healing according to $\Omega$ (i.e., the $\theta$ normalized by its steady state value $\theta_{\rm SS}$), instead of the composite law’s switching according to the slip rate $V$, we could avoid violating the SS canon eq. (26), which the composite law violates around $V=V_{\rm c}$. This mechanism applies to any evolution law conforming eqs. (80) and (83), not limited to the particular examples of eqs. (95) and (96).
@@
Figure 11: Strength evolution predicted by the modified composite law. $A=B$, $L=D_{\rm c}$, $\theta\gg\theta_{\rm X}$, and $\beta=0.001$ are assumed. (a) Predicted strength evolution for ideal VS tests. The visualizing scheme follows that of Fig. 7. Differing from the composite law, the prediction (solid curves) depends only on the ratio $V_{\rm after}/V_{\rm before}$, not not affected by their absolute values because the modified composite law does not have a characteristic speed. Dotted curves show the canonical behaviors (eq. 32), which are not visible for cases with $V_{\rm after}\geq 0.01V_{\rm before}$ as they almost coincide with the prediction by the modified composite law. The inset shows the blowup of the early stage of the evolution for $V_{\rm after}/V_{\rm before}=10^{-3}$, $10^{-4}$. (b) Predicted strength evolution for constant-$\tau$ SHS tests. Following a steady-state sliding at $V_{\rm prior}$, shear stress is abruptly reduced by $\Delta\tau=9A$ and kept constant. The visualizing scheme follows that of Fig. 5. The red wedge indicates the slope of the $\log t_{\rm h}$ asymptote.
@@
The above-mentioned aging-driven rapid strengthening in the earliest part of the transient evolution following a large negative step is a gross violation of the VS canon. On the other hand, Fig. 11a indicates that the subsequent part of the predicted evolution traces the canonical VS behavior following a negative step from $V^{\prime}_{\rm before}$ such that $V^{\prime}_{\rm before}\sim V_{\rm after}/\beta$.
In other words, excepting the very early stage where rapid, time-driven strengthening dominates the prediction, the modified composite law predicts a slip-driven exponential evolution with the fixed characteristic distance specified by the law’s parameter $D_{\rm c}$. This transition is because the modified composite law mutes $f_{t}$ according to $\Omega$; hence, the slip-law-ike behavior takes over in the later stage where $\Omega$ exceeds $\beta$. Contrastingly, the aging law does not have a mute mechanism of $f_{t}$, and the composite law mutes $f_{t}$ according to the absolute level of $V$. Therefore, these two laws do not predict slip-law-like behaviors even later in the transient evolution, and the steady state following a negative step is achieved in a much shorter distance than the parameter $D_{\rm c}$. In this regard, it might be arguable that the modified composite law is still better because it at least reproduces the independence of the characteristic evolution distance $D_{\rm c}$ from the sign and magnitude of the step (Ruina, 1980, 1983). In any case, we emphasize that this switch of the evolution from the log-time healing to the exponential slip relaxation, predicted by the modified composite law, is a direct logical consequence of the general character of our idea eq. (83), or more widely, from eq. (86).
Figure 11b shows the prediction on constant-$\tau$ SHS tests, where shear stress, initialized to the steady state value at $V_{\rm prior}$, is abruptly dropped by $\Delta\tau$ at the beginning of hold, and is then held constant at $\tau_{\rm hold}$ during the hold (§3.3.1). The strength $\Phi$ during the hold is computed using the modified composite law eq. (96). $V$ equals $V_{\rm prior}\exp(-\Delta\tau/A)$ immediately after the beginning of the hold $t_{\rm h}=+0$, and then gradually decreases according to the constitutive law eq. (8) as the strength $\Phi$ increases. During the hold, from eq. (1) with $\theta(0)=D_{\rm c}/V_{\rm prior}$, we see
@@
For interfaces with $A<B$ (Fig. A.1, d3), $\Omega$ decreases as $\theta$ increases, as implied by eq. (97). Hence, if $\Omega\ll\beta$ holds at $t_{\rm h}=+0$, the modified composite law, which satisfies eq. (83), predicts log-time healing throughout $t_{\rm h}>0$, again conforming to the SHS canon. However, for $A>B$ (Fig. A.1, d1), the modified composite law predicts deviation from the canon; although the canonical log-time healing is predicted in the early stage, strengthening will slow down at the time when $\Omega$ comes close to $\beta$ because $\Omega$ increases with $\theta$ on interfaces with $A>B$ (see eq. 97). This canon-violating behavior derives from eq. (83), where $f$ drastically drops from $f\simeq f_{t|{\rm NSC}}=1$ to $f\simeq f_{\delta}V=-(L/D_{\rm c})\Omega\ln\Omega$ (see Appendix A for details). Therefore, this is another side effect of the modified composite law, again a direct consequence of the aging-slip crossover that has brought the aforementioned side effect in large negative VS tests. As far as $L\sim D_{\rm c}$ (eq. 70), this side effect on SHS tests is again a direct logical consequence of the general character of our idea (83) or, more generally, from eq. (86) (see §A.5 for details).
@@ 4.4 Supplemental consideration on $L/D_{\rm c}$
$\psi_{\rm cBase}(\Omega)=\begin{cases}1&(\Omega\ll\beta)\\ \mathcal{O}[(L/D_{\rm c})\beta\ln\beta]&(\beta\ll\Omega\leq 1/e)\\ o[\min_{\pm=\pm 1}(L/D_{\rm c})(1\pm\epsilon_{\Omega}^{*})\ln(1\pm\epsilon_{\Omega}^{*})]&(\Omega\geq 1/e)\end{cases}$ (98)
@@
By multiplying eq. (100) by the Heaviside function $H(1-\epsilon^{*}_{\Omega}-\Omega)$, we can compose $\psi_{\rm cBase}$, as we derived eq. (93) in §4.3. See Appendix E for the allowable range of $\epsilon_{\Omega}^{*}$. Also, as in eq. (94), we can drop the Heaviside truncation as long as we do not mind the slight deviation of $C_{\rm SS}$ from unity. One example to satisfy eq. (100) is the rapidly decaying exponential $\psi_{\rm c}$ (eqs. 95 and 96) proposed for $L\sim D_{\rm c}$ in §4.3, which satisfies eq. (100) for quite a wide range of conditions of $\ln(L/D_{\rm c})=\mathcal{O}(1)$E.3).
Thus, the effect of $L/D_{\rm c}$ is minor in deriving evolution laws that meet the experimental requirements. However, it may be worth noting that the decay rate of $\psi_{\rm cBase}$ required by eq. (98) depends on $L/D_{\rm c}$. For $L\ll D_{\rm c}$, the requirement of eq. (98) is qualitatively the same as that for $L\sim D_{\rm c}$, although more rapid decay is required. For $L\lesssim D_{\rm c}$, $\psi_{\rm cBase}$ satisfying eq. (98) needs to mute the purely time-dependent healing to virtually zero in the range $\Omega\gg\beta$:
@@
which is essentially the same requirement as in §4.3.
By contrast, the requirement of eq. (98) becomes very tolerant for $L\gg D_{\rm c}$ case. In addition to the rapidly decaying $\psi_{\rm cBase}$ (eq. 101), which of course works fine in this case as well, we can even use $\psi_{\rm cBase}$ not muting the purely time-dependent healing at all even for $\Omega\gg\beta$ because, in this case, $f_{\delta}V$ is very “loud” for $\Omega\gg\beta$. Namely, for $L\gg D_{\rm c}$, $|f_{t}|\ll|f_{\delta}V|$ can be secured also by a $\psi_{\rm cBase}$ such that
@@
If the exact satisfaction of eq. (83) is desired, it is enough to multiply $\psi_{\rm cBase}$ by the Heaviside function as done in eq. (93). Appendix F shows the allowable range of $\epsilon_{\Omega}^{*}$ for this example. Note, in such cases as eqs. (102) and (103), where the adopted $\psi_{\rm cBase}$ does not involve $\beta$ as an explicit parameter of the function, appropriate values of $\beta$ must be sought before choosing $\epsilon_{\Omega}^{*}$ accordingly. Appendix F also shows a general procedure to construct appropriate $\beta$ for a given $\psi_{\rm cBase}$, including application cases to eqs. (102) and (103). As discussed there, a certain range of values are considered appropriate as $\beta$, but the order of $\beta$ is fixed when we adopt a weakly decaying $\psi_{\rm cBase}$ (eq. 102).
By introducing another state variable $\theta^{\prime}=(D_{\rm c}/L)\theta$, scaled so that $\theta^{\prime}_{\rm SS}=D_{\rm c}/V$ instead of $L/V$, eq. (103) is converted to
@@
where $\Omega^{\prime}:=V\theta^{\prime}/D_{\rm c}$; note $\dot{\theta}^{\prime}=(D_{\rm c}/L)\dot{\theta}$ and $\Omega^{\prime}=\Omega$. Equation (104) can be read as the sum of the slip law and the purely time-dependent healing of the magnitude $D_{\rm c}/L$ (instead of unity).
When $L\sim D_{\rm c}$, eq. (101), which mutes $f_{t}$ to substantially zero in the non-NSC regime, has been deduced as the unique form of constraint on $f_{t}$4.3). By contrast, as shown above, when $L\gg D_{\rm c}$, eq. (102) ($f_{t}=const.$ except in the close vicinity of steady state) is also acceptable as the constraint on $f_{t}$ alternative to eq. (101). Then, if $L\gg D_{\rm c}$, rather than $L\sim D_{\rm c}$ implicitly presumed in most studies, turns out to be the case, we have to determine, in order to write down evolution laws, which of eq. (101) or (102) is the case. In fact, two evolution laws resulting from eqs. (101) and (102) predict significantly different behaviors, distinguishable in experiments. For example, if eq. (101) is adopted, SHS||||NSC tests on interfaces with $A>B$ are expected to show the slowdown of log-t healing (Fig. A.1, d1) as we detailed for the modified composite law (§4.3; also see §A.5). In contrast, if eq. (102) is adopted, the resulting evolution laws do not predict such slowdown even for interfaces with $A>B$. Thus, provided that the relation between $L$ and $D_{\rm c}$ has not been established, we are ready to write down evolution laws under whatever relation between $L$ and $D_{\rm c}$.
Incidentally, in this §4.4, development for any $L/D_{\rm c}$ has been under the unary ansatz of $\psi_{\rm c}$ (eq. 77, or equivalently 78), and accordingly we sought satisfaction of eq. (83). However, all the arguments in §4.4 hold without this ansatz, by replacing eq. (83) with eq. (86), which applies to any $f_{t}(V,\theta)$, not limited to of $\psi_{\rm c}(\Omega)$ form.

### v3

@@ 4.2.2 Derivation without heuristics
In reading this figure, it must be noted that “aging-/slip-law-like” (i.e. $|f-f_{t}|\lessgtr|f-f_{\delta}V|$) evolution is distinct from evolution “conforming the aging/slip canon” (i.e. $f\simeq f_{t}$, $f\simeq f_{\delta}V$). Rigorous confirmation of the VS-canon conformity of laboratory data is limited to VS tests with $V_{\rm after}/V_{\rm before}>10^{-2}$. VS tests with further lower $V_{\rm after}/V_{\rm before}$ (down to $10^{-3.5}$ so far) are slip-law-like but can significantly differ from the VS canon (Bhattacharya et al., 2015, 2022). Figure 9 leaves this range unpainted. On the other hand, as for SHS, rigorous studies do not exist, to our knowledge, on how closely the laboratory data obeys the aging canon. There is even a possibility that (some of) the results by Nakatani & Mochizuki (1996) are just aging-law-like, not canon-conforming. Under those circumstances, while we painted the whole $(V,\theta)$ space of Nakatani & Mochizuki (1996) down to $\Delta\tau/A=8$ as “canon-conforming,” the formal conclusion of the present paper remains $\beta\lesssim 0.01$, with the lower bound for $\beta$ unspecified (§4.2.1). Nevertheless, finite $\beta$ is instantly concluded mathematically from the SHS tests under $-\log_{10}\Omega\gg 1$ and the VS behavior under $|\log_{10}\Omega|=\mathcal{O}(1)$, not suffering from the current inability to determine the $\beta$ value. Figure 9, though drawn in terms of the concrete values of $\beta$ inferred from experiments probing the currently limited range of $(V,\theta)$ space, is presented only to help readers grasp the general picture of the presently proposed aging-slip switch (i.e., the existence of $\beta$ for $f(V,\theta)$), which has been mathematically proven in §4.2.2 by the deduction from the canons (plus the premise of the TDF).
@@ 4.3 Examples of evolution laws consistent with the three canonical behaviors
In §4.2, we have shown that a class of evolution laws satisfying eqs.(80) and (83) conform constraints from SS, VS, and SHS experiments. The idea is a simple reconciliation of the slip and aging laws (Kato & Tullis, 2001; Ampuero & Rubin, 2008) by switching the two laws according to the value of $\Omega$ ($:=V\theta/L$), the conventional state variable $\theta$ normalized by its steady-state value for the current slip rate $\theta_{\rm SS}(V)$ [$\Omega=\theta/\theta_{\rm SS}(V)$]. The proposed class of laws approximate $\dot{\theta}\simeq 1$ (the aging law at the NSC) for $\Omega\ll\beta$, while they approximate the canonical evolution in VS tests, $\dot{\theta}\simeq-(L/D_{\rm c})\Omega\ln\Omega$ (the slip law). In this subsection §4.3, we propose concrete examples of this class of evolution laws. A stepwise function $\psi_{\rm c}=H(\beta-\Omega)$, where $H(\cdot)$ is the Heaviside function, satisfies eq. (83), of course, but it causes an artificial discontinuity in the evolution function. Instead, we seek a procedure to construct smooth functions satisfying eq. (83). From this subsection onward, we assume $C_{\rm SS}=1$ (eq. 76), whereas §4.2 was developed under a less restrictive assumption of $C_{\rm SS}\sim 1$. In addition, within this subsection §4.3, we impose $L\sim D_{\rm c}$ (eq. 70) to simplify the derivation. As discussed in §3.4, $L\sim D_{\rm c}$ is likely to be the case, if not well confirmed by a wide range of experiments.
Constructing $\psi_{\rm c}$ for $\Omega\gg\beta$ so that it satisfies the first line of eq. (83) is not as straightforward as it might sound. The difficulty comes from $|f_{t|{\rm VS}}|\ll|f_{\delta|{\rm VS}}V|$ (eq. 44) because $f_{\delta}V$ ($=-(L/D_{\rm c})\Omega\ln\Omega$) can become infinitesimally small in the vicinity of $\Omega=1$. To get around, we first make a $\psi_{\rm c}$ that conforms eq. (83) except in the vicinity of $\Omega=1$. Exception handling is next made so that eq. (83) is satisfied also in the vicinity of $\Omega=1$. For this purpose, we introduce here a positive constant $\epsilon_{\Omega}^{*}$ and divide the $\Omega$ range into $|\Omega-1|<\epsilon_{\Omega}^{*}$ and the rest (Fig. 10). To simplify inequality evaluations in the rest of the derivation, we assume $1-\epsilon_{\Omega}^{*}>1/e$, namely $\epsilon_{\Omega}^{*}<1-1/e=0.6321...$, noting that $\Omega\ln\Omega$ takes the unique extremum value at $\Omega=1/e$ (Fig. 10). Also, we hereafter assume $\beta<1/e^{2}(=0.135...)$ and $\sqrt{\beta}\ll 1$, again for simplicity of derivation. As discussed in §4.2, observed symmetry in many existing VS tests has already rejected the possibility that $\beta\geq 1/e^{2}$. Moreover, recent VS tests employing very large steps by Bhattacharya et al. (2015) show that $\sqrt{\beta}\lesssim 0.1$ (eq. 85).
@@
which follows eq. (80) and eq. (76) ($\theta_{\rm SS}=L/V$). In the first place, eq. (34) derived from the VS canon intends that the slip-driven evolution alone makes $\Omega$ approach exactly to its stationary point $\Omega=1$. Hence, involvement of any purely time-dependent healing within the neighborhood of $\Omega=1$, even minuscule, would shift the steady state from $\Omega=1$, meaning that $C_{\rm SS}$ would deviate from unity. Therefore, $\psi_{\rm c}$ needs to be truncated for $\Omega$ almost at the steady state. A straightforward way to realize eq. (91) is to impose $\psi_{\rm c}=0$ in $\Omega>1-\epsilon_{\Omega}^{*}$:
$\psi_{\rm c}(\Omega)=\psi_{\rm cBase}(\Omega)H(1-\epsilon^{*}_{\Omega}-\Omega).$ (93)
@@
Hereafter, we refer to eq. (96) as the ‘modified composite law’ because its form under $L=D_{\rm c}$ is exactly the composite law (eq. 74) by Kato & Tullis (2001), with the constant $V_{\rm c}$ in eq. (74) replaced by $\beta L/\theta$. To sum up, by turning on/off the purely time-dependent healing according to $\Omega$ (i.e., the $\theta$ normalized by its steady-state value $\theta_{\rm SS}$), instead of the composite law’s switching according to the slip rate $V$, we could avoid violating the SS canon eq. (26), which the composite law violates around $V=V_{\rm c}$. This mechanism applies to any evolution law conforming eqs. (80) and (83), not limited to the particular examples of eqs. (95) and (96).
@@
Figure 11: Strength evolution predicted by the modified composite law. $A=B$, $L=D_{\rm c}$, $\theta\gg\theta_{\rm X}$, and $\beta=0.001$ are assumed. (a) Predicted strength evolution for ideal VS tests. The visualizing scheme follows that of Fig. 7. Differing from the composite law, the prediction (solid curves) depends only on the ratio $V_{\rm after}/V_{\rm before}$, not affected by their absolute values because the modified composite law does not have a characteristic speed. Dotted curves show the canonical behaviors (eq. 32), which are not visible for cases with $V_{\rm after}\geq 0.01V_{\rm before}$ as they almost coincide with the prediction by the modified composite law. The inset shows the blowup of the early stage of the evolution for $V_{\rm after}/V_{\rm before}=10^{-3}$, $10^{-4}$. (b) Predicted strength evolution for constant-$\tau$ SHS tests. Following a steady-state sliding at $V_{\rm prior}$, shear stress is abruptly reduced by $\Delta\tau=9A$ and kept constant. The visualizing scheme follows that of Fig. 5. The red wedge indicates the slope of the $\log t_{\rm h}$ asymptote.
@@
The above-mentioned aging-driven rapid strengthening in the earliest part of the transient evolution following a large negative step is a gross violation of the VS canon. On the other hand, Fig. 11a indicates that the subsequent part of the predicted evolution traces the canonical VS behavior following a negative step from $V^{\prime}_{\rm before}$ such that $V^{\prime}_{\rm before}\sim V_{\rm after}/\beta$. In other words, excepting the very early stage where rapid, time-driven strengthening dominates the prediction, the modified composite law predicts a slip-driven exponential evolution with the fixed characteristic distance specified by the law’s parameter $D_{\rm c}$. This transition is because the modified composite law mutes $f_{t}$ according to $\Omega$; hence, the slip-law-ike behavior takes over in the later stage where $\Omega$ exceeds $\beta$. Contrastingly, the aging law does not have a mute mechanism of $f_{t}$, and the composite law mutes $f_{t}$ according to the absolute level of $V$. Therefore, these two laws do not predict slip-law-like behaviors even later in the transient evolution, and the steady state following a negative step is achieved in a much shorter distance than the parameter $D_{\rm c}$. In this regard, it might be arguable that the modified composite law is still better because it at least reproduces the independence of the characteristic evolution distance $D_{\rm c}$ from the sign and magnitude of the step (Ruina, 1980, 1983). In any case, we emphasize that this switch of the evolution from the log-time healing to the exponential slip relaxation, predicted by the modified composite law, is a direct logical consequence of the general character of our idea eq. (83), or the generalized version, eq. (86).
Figure 11b shows the prediction on constant-$\tau$ SHS tests, where, following the initialization by the steady-state sliding at $V_{\rm prior}$, shear stress is abruptly dropped by $\Delta\tau$ at the beginning of hold and is then held constant at $\tau_{\rm hold}$ during the hold (§3.3.1). The strength $\Phi$ during the hold is computed using the modified composite law eq. (96). $V$ equals $V_{\rm prior}\exp(-\Delta\tau/A)$ immediately after the beginning of the hold $t_{\rm h}=+0$, and then gradually decreases according to the constitutive law eq. (8) as the strength $\Phi$ increases. During the hold, from eq. (1) with $\theta(0)=D_{\rm c}/V_{\rm prior}$, we see
@@
For interfaces with $A<B$ (Fig. A.1, d3), $\Omega$ decreases as $\theta$ increases, as implied by eq. (97). Hence, if $\Omega\ll\beta$ holds at $t_{\rm h}=+0$, the modified composite law, which satisfies eq. (83), predicts log-time healing throughout $t_{\rm h}>0$, again conforming to the SHS canon. However, for $A>B$ (Fig. A.1, d1), the modified composite law predicts deviation from the canon; although the canonical log-time healing is predicted in the early stage, strengthening will slow down at the time when $\Omega$ comes close to $\beta$ because $\Omega$ increases with $\theta$ on interfaces with $A>B$ (see eq. 97). This canon-violating behavior derives from eq. (83), where $f$ drastically drops from $f\simeq f_{t|{\rm NSC}}=1$ to $f\simeq f_{\delta}V=-(L/D_{\rm c})\Omega\ln\Omega$ (see Appendix A for details). Therefore, this is another side effect of the modified composite law, again a direct consequence of the aging-slip switch that has brought the aforementioned side effect in large negative VS tests. As far as $L\sim D_{\rm c}$ (eq. 70), this side effect on SHS tests is again a direct logical consequence of the general character of our idea (83) or, more generally, from eq. (86) (see §A.5 for details).
@@ 4.4 Supplemental consideration on $L/D_{\rm c}$
$\psi_{\rm cBase}(\Omega)=\begin{cases}1&(\Omega\ll\beta)\\ \mathcal{O}[(L/D_{\rm c})\beta\ln\beta]&(\beta\ll\Omega\leq 1/e)\\ o[\min_{\pm}(L/D_{\rm c})(1\pm\epsilon_{\Omega}^{*})\ln(1\pm\epsilon_{\Omega}^{*})]&(\Omega\geq 1/e)\end{cases}$ (98)
@@
By multiplying eq. (100) by the Heaviside function $H(1-\epsilon^{*}_{\Omega}-\Omega)$, we can compose $\psi_{\rm c}$, as we derived eq. (93) in §4.3. See Appendix E for the allowable range of $\epsilon_{\Omega}^{*}$. Also, as in eq. (94), we can drop the Heaviside truncation as long as we do not mind the slight deviation of $C_{\rm SS}$ from unity. One example to satisfy eq. (100) is the rapidly decaying exponential $\psi_{\rm c}$ (eqs. 95 and 96) proposed for $L\sim D_{\rm c}$ in §4.3, which satisfies eq. (100) for quite a wide range of conditions of $\ln(L/D_{\rm c})=\mathcal{O}(1)$E.3).
Thus, the effect of $L/D_{\rm c}$ is minor in deriving evolution laws that meet the experimental requirements. However, it may be worth noting that the decay rate of $\psi_{\rm cBase}$ required by eq. (98) depends on $L/D_{\rm c}$. For $L\ll D_{\rm c}$, the constraint of eq. (98) is qualitatively the same as that for $L\sim D_{\rm c}$ although more rapid decay is required. For $L\lesssim D_{\rm c}$, $\psi_{\rm cBase}$ satisfying eq. (98) needs to mute the purely time-dependent healing to virtually zero in the range $\Omega\gg\beta$:
@@
which is essentially the same constraint on $\psi_{\rm cBase}$ as in §4.3.
By contrast, the constraint of eq. (98) becomes very tolerant for $L\gg D_{\rm c}$. In addition to the rapidly decaying $\psi_{\rm cBase}$ (eq. 101), which of course works fine in this case as well, we can even use $\psi_{\rm cBase}$ not muting the purely time-dependent healing at all even for $\Omega\gg\beta$ because, in this case, $f_{\delta}V$ is very “loud” for $\Omega\gg\beta$. Namely, for $L\gg D_{\rm c}$, $|f_{t}|\ll|f_{\delta}V|$ can be secured also by a $\psi_{\rm cBase}$ such that
@@
If the exact satisfaction of eq. (83) is desired, it is enough to multiply $\psi_{\rm cBase}$ by the Heaviside function as done in eq. (93). Appendix F shows the allowable range of $\epsilon_{\Omega}^{*}$ for this example. In such cases as eqs. (102) and (103), where the adopted $\psi_{\rm cBase}$ does not involve $\beta$ as an explicit parameter of the function, appropriate values of $\beta$ must be sought before choosing $\epsilon_{\Omega}^{*}$ accordingly. Appendix F also shows a general procedure to construct appropriate $\beta$ for a given $\psi_{\rm cBase}$, including application cases to eqs. (102) and (103). As discussed there, a certain range of values is considered appropriate as $\beta$, but the order of $\beta$ is fixed when we adopt a weakly decaying $\psi_{\rm cBase}$ (eq. 102).
By introducing another state variable $\theta^{\prime}:=(D_{\rm c}/L)\theta$, scaled so that $\theta^{\prime}_{\rm SS}=D_{\rm c}/V$ instead of $L/V$, eq. (103) is converted to
@@
where $\Omega^{\prime}:=V\theta^{\prime}/D_{\rm c}$. Note $\dot{\theta}^{\prime}=(D_{\rm c}/L)\dot{\theta}$ and $\Omega^{\prime}=\Omega$. Equation (104) can be read as the sum of the slip law and the purely time-dependent healing of the magnitude $D_{\rm c}/L$ (instead of unity).
When $L\sim D_{\rm c}$, eq. (101), which mutes $f_{t}$ to substantially zero in the non-NSC regime, has been deduced as the unique form of constraint on $f_{t}$4.3). By contrast, as shown above, when $L\gg D_{\rm c}$, eq. (102) ($f_{t}=const.$ except in the close vicinity of steady state) is also acceptable as the constraint on $f_{t}$ alternative to eq. (101). Then, if $L\gg D_{\rm c}$, rather than $L\sim D_{\rm c}$ implicitly presumed in most studies, turns out to be the case, we have to determine, in order to write down evolution laws, which of eq. (101) or (102) is the case. Two evolution laws resulting from eqs. (101) and (102) predict significantly different behaviors, distinguishable in experiments. For example, if eq. (101) is adopted, SHS||||NSC tests on interfaces with $A>B$ are expected to show the slowdown of log-t healing (Fig. A.1, d1) as we detailed for the modified composite law (§4.3; also see §A.5). In contrast, if eq. (102) is adopted, the resulting evolution laws do not predict such a slowdown even for interfaces with $A>B$. Thus, provided that the relation between $L$ and $D_{\rm c}$ has not been established, we are ready to write down evolution laws under whatever relation between $L$ and $D_{\rm c}$.
Incidentally, in this §4.4, development for any $L/D_{\rm c}$ has been under the unary ansatz of $\psi_{\rm c}$ (eq. 77, or equivalently, 78), and accordingly, we sought satisfaction of eq. (83). However, all the arguments in §4.4 hold without this ansatz by replacing eq. (83) with eq. (86), which applies to any $f_{t}(V,\theta)$, not limited to of $\psi_{\rm c}(\Omega)$ form.